OpenStudy (18jonea):

If the ratio of the radii of two spheres is 1 : 8, what is the ratio of the surface areas of the two spheres? 1 : 64 1 : 8 1 : 512 r2pi : 8r2pi @michele_laino

OpenStudy (anonymous):

surface area is \[4 \pi r ^{2}\]

OpenStudy (anonymous):

just focus on the radius in the equation

OpenStudy (anonymous):

how does squaring the radius affect both values?

OpenStudy (18jonea):

?

OpenStudy (anonymous):

\[1^{2} : 8^{2}\]

OpenStudy (anonymous):

what does that produce?

OpenStudy (anonymous):

what is \[1^{2}\]

OpenStudy (18jonea):

1

OpenStudy (18jonea):

@10115658

OpenStudy (anonymous):

yep, and what's \[8^{2}\]

OpenStudy (18jonea):

64

OpenStudy (anonymous):

so the answer must be?

OpenStudy (18jonea):

well dont you have to have pi in theresomewhere

OpenStudy (anonymous):

the pi is a constant term so it just cancels out

OpenStudy (18jonea):

so wouldnt it be d or no

OpenStudy (18jonea):

really?

OpenStudy (anonymous):

yep

OpenStudy (18jonea):

If the ratio of the radii of two spheres is 6 : 5, what is the ratio of the surface areas of the two spheres? 36 : 25 6r2pi : 5r2pi 6 : 5 216 : 125 so would this be 36:25

OpenStudy (anonymous):

\[4\pi (1)^{2}=4\pi (8)^{2}\]

OpenStudy (anonymous):

do you see why the 4pi will cancel?

OpenStudy (18jonea):

yes cause both or each sid

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

36:25 is also right

OpenStudy (18jonea):

Thank You!!!!!

OpenStudy (anonymous):

no problem